### v3

@@ 8.3 Experiment 3: learning digenic functions
Materials. We focused on re-discovering a target isoenzyme function associated with the gene $tyrB$ in the Tryptophan biosynthesis pathways. We aim to demonstrate gene-isoenzyme function learning by recovering this gene function. Associated functions of $tyrB$ and $aspC$ can be classified as isoenzyme functions since they are responsible for producing an important amino acid L-phenylalanine. In our logical representation of iML1515, we use reaction and $codes/2$ unit ground literals to describe an isoenzyme. We selected 33 genes associated with this essential pathway. In this focus gene set, 27 genes are associated with a single function and 6 genes are related to two functions. We consider medium conditions with 7 optional nutrients including aromatic amino acids for recovering the deleted $tyrB$ isoenzyme function. Aromatic amino acids are fundamental for protein synthesis but also serve as precursors for vital secondary metabolites and high-value chemicals choi_violacein_2015 .
@@
Methods. We removed tyrB’s gene function and metabolic reaction associated with this isoenzyme from $BMLP_{active}$. The experiment instance space contained double gene-knockout and single gene-knockout experiments of 33 genes. There were $\binom{33}{2}$ gene pair combinations for double-gene knockout experiments and 33 single-gene knockout experiments. We considered growth medium conditions with the 7 optional nutrients. Therefore, the experiment instance space was $(\binom{33}{2}+33)\times 7=3927$ and 3696 experiment instance labels were synthetically generated from double-knockout experiments using the full iML1515 background knowledge. The rest 231 single-knockout empirical data were from validation data in iML1515 .
@@
In the metabolic network, an isoenzyme function hypothesis contains a node represented by a new enzyme identifier (a $enzyme/6$ clause) and a new link between a gene locus identifier and this new enzyme identifier (a $codes/2$ clause). The hypothesis space contained associations of 33 genes (from the focused library) with 33 $enzymes/6$ clauses. There were $(27\times 32+6\times 31)=1050$ potential hypotheses for 27 single-function genes and 6 double-function genes. We also included the tyrB function hypothesis and an empty hypothesis, which gave 1052 hypotheses in total. We increased linearly and exponentially the number of experiments $N$ that could be selected according to {1, 2, 4, 5, 8, 10, 15, 16, 20, 25, 30, 32, 64, 128, 256}. $BMLP_{active}$ selected $N$ experiments from this instance space to actively learn from the hypothesis space. The random selection strategy randomly sampled $N$ instances from the instance space. Again, both methods output the hypothesis with the highest compression according to the sampled examples. A hypothesis was randomly selected when multiple competing hypotheses shared the highest compression. We recovered the models by adding the final hypotheses. We recorded the number of experiments selected and evaluated if the recovered models contained the target isoenzyme function from the iML1515 background knowledge. Each experiment selection method was repeated 10 times and we computed the frequency of successful isoenzyme recovery across these 10 repeats. We show this comparison in Fig 8.
Results and discussion. Fig 8 shows the result of recovering tyrB’s isoenzyme function. For answering question Q5, we observed that $BMLP_{active}$ successfully recovers the correct tyrB isoenzyme function in 20 experiments. This demonstrates $BMLP_{active}$’s ability to discover gene-isoenzyme associations in GEM. However, to learn all isoenzyme functions, a high-throughput experimentation procedure is required and we further discuss this in Section 9.
This result also shows that $BMLP_{active}$ significantly reduced the number of experiments needed compared to random experiment selection, which answers the question Q3. $BMLP_{active}$ provides higher information gain from each experiment and can guarantee recovery with as few as 20 experiments. While random sampling could prune most candidate hypotheses, it failed to eliminate competitive hypotheses further. In contrast to $BMLP_{active}$, random sampling could not recover this isoenzyme function with 64 experiments and could only do this occasionally with more than 10 times the number of experiments required by $BMLP_{active}$. Although we used a reduced set of genes in this demonstration, the combinatorial space was sufficiently large that random experiment selection became non-viable as an experimentation strategy in a discovery process. The remarkable increase in efficiency with $BMLP_{active}$ with this task demonstrates it has potential even as the experimental design space grows.
9 Conclusion and future work
We present a new logic programming system $BMLP_{active}$ that can simulate and learn gene functions in a state-of-the-art GEM. This GEM, iML1515, describes 1515 genes and 2719 metabolic reactions in E. coli. To our knowledge, $BMLP_{active}$ is the first logic programming system able to operate on GEMs. We implement $BMLP_{active}$ via our proposed BMLP approach by exploiting boolean matrices to improve the simulation runtime of a GEM by 170 times. Our empirical results show that $BMLP_{active}$ can learn gene functions while reducing the cost of optional nutrient substance in experiments by 90%. In addition, we deploy $BMLP_{active}$ to actively learn gene-isoenzyme functions which are an error source in GEM. As a demonstration, we empirically show that $BMLP_{active}$ can learn an important isoenzyme function in the aromatic amino acid biosynthesis pathway with as little as 20 training examples. In comparison, random experiment selection cannot learn this isoenzyme function consistently even with more than 250 training data. $BMLP_{active}$ provides a substantial advantage for isoenzyme function discovery over random experiment selection due to the sample complexity reduction. This finding suggests that $BMLP_{active}$ could be used to offset the growth in experimental design spaces associated with genome-scale function learning.
One limitation of $BMLP_{active}$ is that it cannot yet quantitatively predict the behaviours of biological systems. Quantitative growth predictions would allow us to model in-vivo phenotype observations without binary discretisation. Future work will focus on using logic programs to optimise relations and parameters given numeric training data. A natural link exists between Petri nets and Probabilistic Logic Programming de_raedt_probabilistic_2015 . Transitions sometimes have firing likelihood constraints and the uncertainty can be modelled by probabilistic logic programs. Evaluating these programs requires tensor-based algebra of logic programs, a topic that has been explored recently in sakama_logic_2021 .
$BMLP_{active}$ also requires a pre-defined hypothesis space since it does not invent predicates for learning new concepts. This relates to a central problem of ILP known as predicate invention ILP30 . New predicate names and reusable definitions automatically learned in addition to the background knowledge are machine inventions. Recent methods have combined ILP with meta-level logic programs and matrices such as the Deeplog system muggleton_hypothesizing_2023 , $\delta$ILP evans_learning_2018 and $Meta_{Abd}$ dai_abductive_2021 . Reusing invented predicates can reduce the size of a hypothesis learned by an ILP system and improve its learning performance ILP30 . $BMLP_{active}$ could be extended with predicate invention for learning novel biological knowledge.
To make a biological discovery, we aim to integrate this approach with an experiment workflow such as the one shown in wang_pooled_2018 to address whole-genome function learning. Data created from a high-throughput empirical framework can be first compared with simulated phenotypic predictions. The inconsistency can inform us of the error sources in the biological model, narrowing the scope of hypothesis space. With advanced biological gene editing tools, genes edition can be targeted for experimental confirmation of candidate hypotheses during active learning. Importantly, we show that our approach can address the interactions between multiple genes, which are not addressable by current mutagenesis approaches.
Designing experiments to explain some hypothesised phenomenon requires deliberate thinking by scientists. A helpful scientific assistant could teach scientists to create experiment plans with optimal outcomes. This aspect of human-AI collaboration has been studied in an emerging research area on Ultra-Strong Machine Learning (USML) michie1988 . A USML not only learns a concept but also teaches this concept to humans such that human comprehension is improved. Prior research has shown instances of USML in the context of learning recursive concepts US2018 and a two-player game Ai2021 . A recent study sequential_teaching demonstrated that an ILP system can play the role of a “teacher” and human learning from machine explanations of its output can sometimes inspire human discovery. To date, there is no demonstration of autonomous discovery systems that satisfy the USML criterion. The ultimate goal of autonomous discovery systems is to acquire new scientific understanding and explain insights to humans krenn_scientific_2022 . The present work can be extended to examine the effect of machine-selected experiments and machine-learned output on human experiment selection and human comprehension of complex biological concepts.
Ethical Statement
Datasets used do not contain any personal information or other sensitive content. There are no ethical issues.
Data Availability
Data used and generated in all experimental sections are accessible on
https://github.com/lai1997/BMLP_active_public.
Acknowledgments
Appendix A. Reagent (optional nutrient substance) costs
@@
Appendix B. Gene function learning example distribution in Experiment 2.
Table 4: The distribution of synthetic positive and negative examples for each gene knockout based on simulation of iML1515. Gene identifiers start with the lower-case letter “b” and gene descriptive names are given in parentheses.

### v4

@@ G.4 Experiment 3: learning digenic functions
Materials. We focused on re-discovering a target isoenzyme function associated with the gene $tyrB$ in the Tryptophan biosynthesis pathways. We aim to demonstrate gene-isoenzyme function learning by recovering this gene function. Associated functions of $tyrB$ and $aspC$ can be classified as isoenzyme functions since they are responsible for producing an important amino acid L-phenylalanine. In our logical representation of iML1515, we use reaction and $codes/2$ unit ground literals to describe an isoenzyme. We selected 33 genes associated with this essential pathway. In this focus gene set, 27 genes are associated with a single function and 6 genes are related to two functions. We consider medium conditions with 7 optional nutrients, including aromatic amino acids for recovering the deleted $tyrB$ isoenzyme function. Aromatic amino acids are fundamental for protein synthesis but also serve as precursors for vital secondary metabolites and high-value chemicals [13].
@@
Methods. We removed tyrB’s gene function and metabolic reaction associated with this isoenzyme from $BMLP_{active}$. The experiment instance space contained double gene-knockout and single gene-knockout experiments of 33 genes. There were $\binom{33}{2}$ gene pair combinations for double-gene knockout experiments and 33 single-gene knockout experiments. We considered growth medium conditions with the 7 optional nutrients. Therefore, the experiment instance space was $(\binom{33}{2}+33)\times 7=3927$ and 3696 experiment instance labels were synthetically generated from double-knockout experiments using the full iML1515 background knowledge. The rest 231 single-knockout empirical data were from validation data in [43].
@@
In the metabolic network, an isoenzyme function hypothesis contains a node represented by a new enzyme identifier (a $enzyme/6$ clause) and a new link between a gene locus identifier and this new enzyme identifier (a $codes/2$ clause). The hypothesis space contained associations of 33 genes (from the focused library) with 33 $enzymes/6$ clauses. There were $(27\times 32+6\times 31)=1050$ potential hypotheses for 27 single-function genes and 6 double-function genes. We also included the tyrB function hypothesis and an empty hypothesis, which gave 1052 hypotheses in total. We increased linearly and exponentially the number of experiments $N$ that could be selected according to {1, 2, 4, 5, 8, 10, 15, 16, 20, 25, 30, 32, 64, 128, 256}. $BMLP_{active}$ selected $N$ experiments from this instance space to actively learn from the hypothesis space. The random selection strategy randomly sampled $N$ instances from the instance space. Again, both methods output the hypothesis with the highest compression according to the sampled examples. A hypothesis was randomly selected when multiple competing hypotheses shared the highest compression. We recovered the models by adding the final hypotheses. We recorded the number of experiments selected and evaluated if the recovered models contained the target isoenzyme function from the iML1515 background knowledge. Each experiment selection method was repeated 10 times, and we computed the frequency of successful isoenzyme recovery across these 10 repeats.
Appendix H Reagent costs of optional nutrient substances
@@
Appendix I Gene function learning example distribution in Experiment 2.
Table 4: The distribution of synthetic positive and negative examples for each gene knockout based on iML1515. Gene identifiers start with the lower-case letter “b” and gene descriptive names are given in parentheses.